the height of a radio tower is 450 feet. how long should a guy wire be if it is to connect to the middle of the tower and be secured at point 100 feet from the base?
Found 2 solutions by mananth, nyc_function:
Answer by mananth(16949)

tower height = 450 feet
half of 450 = 225
height at which wire is = 225 feet
distance on ground 100 feet.
..
Apply pythagoras theorem
wire is the hypotenuse.
wire length = sqrt ( 225^2 +100^2)
= 246.22

Answer by nyc_function(2741)

Use the Pythagorean Theorem.
Let G = guy wire length
(225)^2 + (100)^2 = G^2
NOTE: The guy wire is connected 1/2 of 450 feet above the ground and it also extends to a point 100 feet away from the base of the tower. This means the guy wire is connected at a height of 225 feet.
Can you finish?